OpenStudy (anonymous):

Suppose the ice cream on top of a cone makes a perfectly spherical shape. If the diameter of the ice cream scoop was 6 inches, what will happen to the volume of the ice cream when the ice cream scoop is decreased to a diameter of 3 inches? What will the original and new volume of the ice cream be?

OpenStudy (anonymous):

the cone does not change the ice cream does.

OpenStudy (anonymous):

not at all....

OpenStudy (anonymous):

ok one sec let me explain.

OpenStudy (anonymous):

Volume = Length * width * height v= l*w*h v=6*6*3 they are all the same around the whole shape , think of it as a circle next is the new volume which is taking half of the original off, we are decreasing the circle Volume = Length * width * height v=l*w*h v=3*3*1.5 Diameter is the outside of the circle with the dimensions of 6

OpenStudy (anonymous):

so does that make 6 the answer?

OpenStudy (anonymous):

& how do I see what you said earlier? Its not showing anymore

OpenStudy (anonymous):

I dont know what happened but no its multiplication then subtraction 3x3x1.5-6 6x6x3

OpenStudy (anonymous):

thank you! I get it now

OpenStudy (anonymous):

your welcome :)